A hollow sphere starts from rest and rolls down a hill without sliding. At the bottom of the hill, it has a linear velocity of 5 m/s. What was the height of the hill the sphere rolled down (in meters)?
Answer #1

from the law of conservation on energy

mgh = 0.5 mv^2 + 0.5 I w^2

I is moment of inertia = 2mr^2/3

W = V/r

so


gh = 0.5 mv^2 + 1/3 * m r^2 * V^2/r^2

9.8 * h = 0.5* 5^2 + 1/3 * 5^2

h = 2.125 m